# Expert-Guided Prompting and Retrieval-Augmented Generation for Emergency Medical Service Question Answering

### v1

@@ A.7.2 Detailed statistics of Benchmarks
We conducted an error analysis on our best model, ExpertRAG-GT + Expert-CoT, which achieved 82.24% accuracy. From the 661 errors on the test set, we randomly selected 50 samples to manually examine the questions, retrieved knowledge, and model reasoning. As shown in Figure 9, we attribute the errors to five main categories:
1) Reasoning Error (42.4%): the retrieved documents are helpful to answer the question and the model used the knowledge but the reasoning is incorrect.
2) Retrieval Error(20.3%): no relevant knowledge is retrieved, and model reasoning by “model knowledge””, leading to the wrong answer.
3) Fail to Use the Correct Knowledge (20.3%): part of the retrieved knowledge is helpful to answer the question but the model use other irrelevant knowledge for reasoning.
4) Misinterpret the Question (15.3%): the model didn’t interpret the question correctly. For example, one question is what is the age cutoff for using an aed on a patient? and the choices are a. 18 year old, b. 5 year old, c. 25 year old, d. 1 year old. The question is asking about minimum age (d. 1 year old) for routine AED use. However the model’s understanding is adult age threshold (a. 18 year old).
5) Model Policy Constraint (1.7%): the model think none of the choices is correct and abstain to answer.

### v2

@@ A.9.1 Error Types
Figure 10: Error Rate per Subject Area
Figure 11: Error Rate per Certification
A.9.2  Errors per Subject Area and Certification
To investigate how errors are distributed across subject areas and certification levels, we compare the error rates of three models, (Qwen3-4B (0-shot), ExpertRAG-4B (Expert-CoT) and Qwen3-32B (0-shot)). The results are shown in Figure 10 and Figure 11. In subject areas, all three models make most of their mistakes in the core NREMT domains: “airway”, “EMS operations”, “medical–OB/GYN”, “trauma”, and “pediatrics”. Together, these categories account for the majority of errors, while “anatomy”, “assessment”, “pharmacology”, and “others” contribute relatively few. Across certification levels, all models exhibit the highest error rates at the EMT level, followed by Paramedic and EMR, with AEMT contributing the fewest errors.
ExpertRAG-4B consistently reduces the absolute number of errors compared to the Qwen3-4B baseline across nearly all subject areas and certification levels. The Qwen3-32B has the fewest errors across three models. One finding is although the total number of errors are reduced in ExpertRAG-4B and Qwen3-32B models, the relative distribution of errors across subject areas is still the same for each model.